### v1

# HAIC: Improving Human Action Understanding and Generation with Better Captions for Multi-modal Large Language Models

Abstract
Recent Multi-modal Large Language Models (MLLMs) have made great progress in video understanding. However, their performance on videos involving human actions is still limited by the lack of high-quality data. To address this, we introduce a two-stage data annotation pipeline. First, we design strategies to accumulate videos featuring clear human actions from the Internet. Second, videos are annotated in a standardized caption format that uses human attributes to distinguish individuals and chronologically details their actions and interactions. Through this pipeline, we curate two datasets, namely HAICTrain and HAICBench. HAICTrain comprises 126K video-caption pairs generated by Gemini-Pro and verified for training purposes. Meanwhile, HAICBench includes 500 manually annotated video-caption pairs and 1,400 QA pairs, for a comprehensive evaluation of human action understanding. Experimental results demonstrate that training with HAICTrain not only significantly enhances human understanding abilities across 4 benchmarks, but can also improve text-to-video generation results. Both the HAICTrain and HAICBench are released at https://huggingface.co/datasets/KuaishouHAIC/HAIC.
HAIC: Improving Human Action Understanding and Generation with Better Captions for Multi-modal Large Language Models
Xiao Wang $1$Equal contribution. [^2] Work done as an intern at Kuaishou Technology.  Jingyun Hua $1$Equal contribution.  Weihong Lin $1$Equal contribution.  Yuanxing Zhang 1 Fuzheng Zhang 1Jianlong WuDi Zhang 1Liqiang Nie  [^3] Corresponding author. 1Kuaishou Technology
1 Introduction
Multi-modal large language models have notably showcased their preeminence across various video understanding tasks Chen et al. (2024d); Li et al. (2024a); Wang et al. (2024a); Zhang et al. (2024b). Among these, human action understanding plays a critical role in many downstream applications, e.g., human-computer interaction Hayes (2011), autonomous driving Xu et al. (2021), embodied intelligence Gupta et al. (2021), and human video generation Wang et al. (2024b).
A recent study, ShareGPT4Video Chen et al. (2024a) has demonstrated that high-quality and detailed video captions can improve MLLMs’ performance in video understanding. However, most existing works Soomro et al. (2012); Xu et al. (2016); Krishna et al. (2017); Wang et al. (2019); Chen et al. (2024c) provide only coarse captions for human actions, insufficient for understanding fine-grained behaviors. MotionLLM Chen et al. (2024b) introduces the MoVid dataset with fine-grained action captions from MotionX Lin et al. (2023). Nevertheless, this dataset mainly focused on single-person scenarios. For multi-person situations, MoVid only considered consistent group activities like “a group of people performing the Korean dance”. A comprehensive dataset is essential to enhance MLLMs’ understanding of detailed human actions and interactions in both single- and multi-person contexts, critical for tasks like emotional analysis, motivation prediction, and relationship modeling.
Figure 1: Our standardized caption format presents each individual’s detailed attributes, body actions, and interactions in chronological order, making it easier to distinguish individuals and comprehend their behaviors.
There are two main challenges for building such datasets: (1) Action Video Accumulating. How to automatically accumulate large-scale videos featuring clear actions of multiple individuals. (2) Caption formatting. How to define a caption format that can clearly distinguish different people and detail their behaviors and interactions respectively.
To address the above challenges, we propose a novel data generation pipeline composed of two stages. In the video accumulation stage, we accumulate videos from various domains featuring clear, meaningful human actions and identify their specific timestamps. This process is highly selective, with only about 1% of videos meeting our quality criteria after applying various strategies. In the caption annotation stage, we define a caption format that uses human attributes to distinguish individuals and chronologically annotates detailed body actions and interactions for each person (see Figure 1). With this pipeline, we curate two datasets: HAICTrain (Human Action and Interaction Comprehension Training set) and HAICBench (a benchmark for evaluation). HAICTrain contains 126K videos accumulated from WebVid Bain et al. (2021), annotated in our defined caption format by Gemini-1.5-Pro Team et al. (2024). HAICBench includes 500 YouTube videos with human-annotated captions in the same format. Furthermore, we generate 1,400 multiple-choice QA pairs across five categories—human interaction, action detail, action sequence, count, and human attribute—by prompting GPT-4o Hurst et al. (2024) and Gemini-1.5-Pro. Note that all the above machine annotation results undergo review and refinement by human annotators. Experimental results indicate that utilizing HAICTrain for training can remarkably enhance the model’s human action understanding ability by 1.4%-2.1%. Additionally, in MovieGenBench Polyak et al. (2024), our post-trained model surpasses the original model in GSB score of 2.15 and 6.81 in HunyuanVideo Kong et al. and Wanx2.1 Aliyun (2023), respectively.
Category Example
What gesture does the middle-aged woman make while talking to the other man?
Interaction (A) Gestures with both hands clasped in front of her (B) Claps her hands (C) Waves her hands in the air (D) Points at the desk
What does the man in the black hat do with his right hand as he starts skateboarding?
Action Details (A) He waves it in greeting (B) He points down the slope (C) He places it on his hip (D) He keeps it in his pocket
What does the man in the gray cap do immediately after gripping the barbell?
Action Sequence (A) He looks at the camera (B) He adjusts his hat (C) He bends down and lowers the barbell (D) He walks towards his front left and rubs his hands together
How many times does the man clap his hands?
Count (A) Three (B) Four times (C) Once (D) Twice
What color is the cropped jacket worn by the young female character in the video?
Attribute (A) Pink (B) White (C) Blue (D) Black
Table 1: Task examples of HAICBench, showcasing a comprehensive human action understanding across spatial (action details), temporal (sequence, count), and multi-human interaction (interaction and attribute) aspects.
Our contributions can be summarized as follows:
  • We propose a novel data annotation pipeline to provide data that can facilitate human action understanding, which 1) accumulates large-scale videos with clear actions from the Internet and 2) generates standardized captions that distinguish individuals and detail their actions and interactions.
  • We introduce two datasets: HAICTrain which includes 126K generated-then-verified high-quality video-caption pairs for training; and HAICBench comprising 500 human annotated video-caption pairs and 1,400 QA pairs, designed to evaluate MLLMs’ human action understanding comprehensively.
  • Experiments demonstrate that training with HAICTrain significantly improves human action understanding in benchmarks including MVBench Li et al. (2024c), PerceptionTest Patraucean et al. (2024), ActivityNet-QA Yu et al. (2019) and HAICBench. Furthermore, HAICTrain substantially improves text-to-video generation on MovieGenBench.
Figure 2: Our data generation pipeline. (a) The video accumulation stage collects videos featuring clear human actions from the Internet. Based on this, (b) HAICTrain is curated through Gemini-1.5-Pro re-captioning, and (c) HAICBench is created by LLM-assisted human annotation.
2 Related Work
2.1 Video Caption Datasets
Most existing video captioning datasets prioritized general video understanding Xu et al. (2016); Zhou et al. (2018); Wang et al. (2019); Miech et al. (2019); Bain et al. (2021); Zellers et al. (2021); Lei et al. (2021); Xue et al. (2022); Chen et al. (2024a, c); Wang et al. (2024d, c); Xiong et al. (2024); Wang et al. (2024b). These datasets included only a subset of human videos, and the action captions tend to be coarse. Some datasets specifically focused on human-centric captioning, including ActivityNet-Captions Krishna et al. (2017), Ego4D Grauman et al. (2022), Kinetic-GEB Wang et al. (2022), ActionHub Zhou et al. (2024) and OpenHumanVid Li et al. (2024b).
Different from these datasets, our work focuses on finer-grained human actions and interaction understanding.
2.2 Video Understanding Benchmarks
Traditional video understanding benchmarks have primarily honed in a fixed set of classes Soomro et al. (2012); Shahroudy et al. (2016); Carreira and Zisserman (2017); Goyal et al. (2017). Recently, there has been a shift towards open-set video understanding, including video captioning and question answering Xu et al. (2016); Wang et al. (2019); Xiao et al. (2021); Li et al. (2024c). While benchmarks like NeXT-QA Xiao et al. (2021), MVBench Li et al. (2024c), and PerceptionTest Patraucean et al. (2023) include action-related questions, they primarily focus on general action recognition rather than fine-grained details. TemporalBench Cai et al. (2024) evaluates event-level action understanding using a different protocol that requires distinguishing correct captions from negative ones. Action-specific benchmarks, such as ActivityNet-Captions Krishna et al. (2017), TGIF-QA Jang et al. (2019), MoVid-Bench Chen et al. (2024b), EGOBody Zhang et al. (2022), and GRAB Taheri et al. (2020), address temporal sequences or human-object interactions but often overlook human-human interactions.
Our HAICBench provides a more comprehensive evaluation of human action understanding, covering five key aspects detailed in Table 1.
2.3 Multi-modal Large Language Models
MLLMs generally process multi-modal input information and generate language output. Kosmos Huang et al. (2023) introduced an end-to-end framework that integrated visual inputs with LLM from cohesive training. Flamingo Alayrac et al. (2022) and InstructBLIP Dai et al. (2023) merged visual and linguistic features through cross-attention and a Q-Former module, respectively. MiniGPT-4 Zhu et al. (2024) and LLaVA Liu et al. (2023) simplified the integration by linearly projecting visual features directly into the LLM embedding space.
Recent studies focused on different aspects to enhance the above early attempts in MLLMs. Cosmos-2 Peng et al. (2024) and NeXT-GPT Wu et al. (2024) have expanded MLLM applications to broader multi-modal tasks. LLaVA-1.5 Liu et al. (2024) explored adding high-quality multi-task training data, and scaling up the resolution and LLM size to boost MLLM performance. LLaVA-OneVision Li et al. (2024a) explored to unify dynamic image resolution, multi-image, and video into a unified input format.
3 HAIC Data Pipeline
In this section, we detail our data generation pipeline, as illustrated in Figure 2.
3.1 Action Video Accumulation
This pipeline aims to accumulate human videos featuring clear human actions and sufficient details from large-scale videos.
Metadata Filtering. We begin by discarding low-resolution videos and those without verbs in their descriptions using spaCy Honnibal et al. (2020). The remaining videos are split into short clips with unique scenes using SceneDetect Castellano (2024). We keep clips between 5 and 20s long, as actions typically occur within a single scene.
Human Existence Filtering. We uniformly sample 16 frames from each clip and use the RTMPose Jiang et al. (2023) object detector to identify humans. Only videos where all frames contain 1-5 humans and the total bounding box area covers at least 10% of the frames are retained, ensuring sufficient human detail.
Human Action Filtering. Using RTMPose Jiang et al. (2023), we detect human bounding boxes and 17 body keypoints at 1 fps. Tracklets are constructed based on the maximal IoU between frames. We then filter out videos with static humans by ensuring the $L_{1}$ distance between all adjacent keypoints exceeds 0.085, with keypoint coordinates normalized by the video resolution. This filtering captures clear human actions.
However, we observe that 15% of the filtered videos still contain static humans despite large keypoint $L_{1}$ distances. This often results from camera movements or image gallery videos. In these cases, we empirically find keypoints approximately follow an affine transformation (e.g., translation, scaling, rotation, and shear mapping). Based on this insight, we developed a strategy to filter these videos. Formally, let the keypoint vector in frame $t$ be $\mathbf{P}_{t}\in\mathbb{R}^{3\times 17}$, where $17$ is the number of keypoints and $3$ corresponds to homogeneous coordinates $(\textmd{height},\textmd{width},1)$. We assume keypoints in these videos generally adhere to the affine transformation:
$\begin{cases}\mathbf{P}_{t+1}=\mathbf{T}\mathbf{P}_{t},\\ \mathbf{T}=\begin{bmatrix}\mathbf{A}&\mathbf{t}\\ \mathbf{O}&1\end{bmatrix},\end{cases}$ (1)
where $\mathbf{A}\in\mathbb{R}^{2\times 2}$, $\mathbf{t}\in\mathbb{R}^{2\times 1}$ is the transformation coefficients, and $\mathbf{O}=[0,0]$. We solve the following least squares problems:
$\min_{\mathbf{A},\mathbf{t}}{\left\|\mathbf{P}_{t+1}-\mathbf{T}\mathbf{P}_{t}\right\|_{2}},$ (2)
and retain only those samples with a residual value $r>0.0016$. A larger residual indicates a greater deviation from Equation 1, suggesting that the unwanted videos mentioned above.
Overall, the whole action video accumulation step yields 0.31% to 1.3% of human action videos, depending on the video source.
3.2 HAICTrain
We chose the WebVid-10M dataset Bain et al. (2021) as the video source for training due to its large scale and high vision quality. Initially, we apply the action video accumulation pipeline in Section 3.1 to collect action videos from the WebVid-10M. This process results in a collection of 126K videos, representing 1.2% of the original dataset. Then, we employ Gemini-1.5-Pro Team et al. (2024) to generate captions in the standardized format as depicted in Figure 1 referring to the videos and original captions. The specific prompts used for this process are detailed in Appendix A. We then employ an additional judgement to filter out failure cases that do not follow the pre-defined format or display low quality. This judgement ensures the quality of our data in HAICTrain.
Figure 3: Statistics of HAICBench. Although videos are relatively short, the video captions are of high details including various action details and sequential actions.
3.3 HAICBench
We develop an LLM-assisted human annotation pipeline to create the HAICBench, which evaluates the capabilities of MLLMs in human action understanding. Human annotators first craft video captions following the format in Figure 1. To enhance question diversity, we then adopt LLMs to generate QA pairs based on these captions, which are then verified by annotators.
Video Caption Annotation. To avoid potential overlap with public benchmarks, we choose the newly proposed Omega-multimodal dataset OMEGA Labs (2024) as our benchmark video source, which comprises over 30 million 2-minute video clips. We apply our video accumulation stage in Section 3.1 to obtain some video clips, and manually check these clips to obtain 500 video clips. Then, human annotators are required to describe the number of human subjects, noting static attributes (gender, age, clothing, accessories) for subject identification, followed by body movements, action sequences, and interactions with others in chronological order.
To ensure the quality of annotated captions, we train all annotators for one week. Besides, each video is first annotated by 1 annotator and then checked and made up missing points by another 3 ones. We do not follow previous works like VATEX Wang et al. (2019) or MSR-VTT Xu et al. (2016) to annotate multiple captions for one video, since our caption format is an exhaustive description for human actions in a video.
QA Pair Generation. Based on captions above, we prompt GPT-4o Hurst et al. (2024) and Gemini-1.5-Pro Team et al. (2024) to generate multiple-choice question-answer pairs about human interactions, action details, action sequences, count, and human attributes. The prompts share a similar format with those in MVBench Li et al. (2024c). See Appendix C for details. Each question-answer pair is checked by 2 annotators and will be corrected if there are any mistakes. All options are shuffled to avoid potential bias.
We get a total of 500 video-caption pairs and 1,400 QA pairs for our HAICBench dataset. Detailed statistics are presented in Figure 3. Although each video clip focuses on a single scene and is relatively short (less than 20 seconds), the captions are highly detailed, often exceeding 100 words. The word cloud analysis reveals that our captions provide comprehensive descriptions of actions.
4 Experiments
Model Frames Avg.
Action
Details
Action
Sequence
Interaction Count Attribute
Human Annotated Captions - 96.7 100.0 94.3 95.0 95.3 99.0
Gemini-1.5-Pro Team et al. (2024) - 41.5 29.1 28.7 36.3 54.7 58.7
GPT-4o Hurst et al. (2024) 50* 40.3 32.6 28.3 38.0 52.0 50.7
VideoLLaMA2-7B Cheng et al. (2024) 64 21.5 11.7 18.0 14.7 39.3 24.0
LongVA-7B Zhang et al. (2024a) 64 16.2 11.4 13.0 11.0 25.3 20.3
InternVL2-8B Chen et al. (2024d) 64 29.1 20.0 18.3 22.7 43.3 41.0
Qwen2VL-7B Wang et al. (2024a) 64 28.7 16.6 21.7 21.3 44.7 39.3
LLaVA-Video-7B Zhang et al. (2024b) 64 30.7 19.1 16.0 25.3 44.0 49.0
LLaVA-Video-ActionPro-7B 64 35.7 22.3 17.0 31.7 53.3 54.3
Table 2: Results on HAICBench in the caption evaluation setting. We adopted MLLMs to generate video captions, based on which we prompted an LLM to answer video questions. *The GPT-4o API supports a maximum of 50 frames per video.
Model Frames Avg.
Action
Details
Action
Sequence
Interaction Count Attribute
Gemini-1.5-Pro Team et al. (2024) - 66.9 60.3 49.3 75.7 66.7 82.7
GPT-4o Hurst et al. (2024) 50* 60.6 59.1 43.4 64.0 54.0 82.3
VideoLLaMA2-7B Cheng et al. (2024) 64 49.7 35.4 37.7 48.3 62.7 64.3
LongVA-7B Zhang et al. (2024a) 64 50.3 43.1 39.3 56.0 37.3 75.7
InternVL2-8B Chen et al. (2024d) 64 58.2 41.4 40.0 61.7 68.0 80.0
Qwen2VL-7B Wang et al. (2024a) 64 64.4 61.1 44.7 65.3 71.3 79.7
LLaVA-Video-7B Zhang et al. (2024b) 64 64.3 53.7 43.7 67.3 70.7 86.3
LLaVA-Video-ActionPro-7B 64 66.4 56.6 45.0 68.7 76.0 86.0
Table 3: Results on HAICBench in the standard evaluation setting. *The GPT-4o API supports a maximum of 50 frames per video.
Base Model Training Dataset MVB PerTest ANetQA HAICBench
LLaVA-Video-178K HAICTrain WebVid
200K 126K 126K 126K
LLaVA-Video 59.4 58.1 61.8 64.3
60.0 58.3 63.8 64.7
62.1 59.9 65.2 66.4
58.8 55.8 62.1 62.2
LLaVA-OneVision 57.8 47.9 57.1 61.6
59.3 55.7 61.8 62.2
60.6 56.9 62.7 64.1
58.8 55.8 60.4 61.8
Table 4: The gain from training with our high-quality human action captions is effective across several benchmarks. MVB, PerTest, and ANetQA denotes MVBench, PerceptionTest, and ActivityNet-QA, respectively.
4.1 Experimental Settings
Baselines. We selected LLaVA-Video-7B Zhang et al. (2024b) as our baseline due to its minimal inductive bias in model architecture and superior performance. Furthermore, we enhanced its action understanding capability through post-training, resulting in LLaVA-Video-ActionPro-7B.
Training dataset. To preserve the general capability of LLaVA-Video against catastrophic forgotten, we randomly selected 200K instruction pairs from its training set LLaVA-Video-178K Zhang et al. (2024b) for sample rehearsal Verwimp et al. (2021). Then, this subset is combined with our HAICTrain to form a training set with 326K instruction pairs in total.
Evaluation dataset. To evaluate human action understanding ability, we executed experiments on several action-related benchmarks, including MVBench, ActivityNet-QA, PerceptionTest, and our HAICBench. Note that we focused on questions related to human actions in these benchmarks. For MVBench, we performed a comparison of all sub-tasks whose names contain “Action”, resulting in 7 types: Action Antonym, Action Count, Action Sequence, Action Prediction, Action Localization, Fine-grained Action, and Unexpected Action. For PerceptionTest, we also selected all questions whose tags are related to action. More details are presented in Appendix B.
Evaluation metrics. For multiple-choice questions in MVBench, PerceptionTest, and HAICBench, we followed the prompting approach in LLaVA-1.5 Liu et al. (2024) and used accuracy as the metric. For ActivityNet-QA which features open-ended questions, we followed the evaluation protocol in Video-ChatGPT Maaz et al. (2024), utilizing GPT-4o-0513 to calculate accuracy. In terms of HAICBench, we implemented two evaluation settings: (1) Standard Evaluation, where the video and question are directly input into an MLLM to generate an answer; and (2) Caption Evaluation, where the model generates a caption for each video, and an LLM (Gemini-1.5-Pro) answers questions based on this caption, thereby assessing caption ability based on QA accuracy. To avoid the LLM guessing answers when captions do not mention related contents, we asked Gemini-1.5-Pro to refuse to answer those questions (see Appendix D).
Implementation Details. We used LLaVA-Video-7B as our initial model and fine-tuned it on the training dataset. All parameters were updated during training. The model was fine-tuned for one epoch with a learning rate of 1e-5 for the LLM and 2e-6 for the vision encoder, with 256 batch size. We sampled 64 frames uniformly in both the training and evaluation stages. During generation, we take the greedy search without randomness. Our experiments utilized 128 NVIDIA A800-80GB GPUs.
4.2 Results on HAICBench
Caption Evaluation Setting. We follow the caption evaluation setting of HAICBench, where MLLM baselines generate action-related captions for each video, and then the captions and questions are fed into Gemini-1.5-Pro for answers. The results are presented in Table 2. Our post-trained LLaVA-Video-ActionPro-7B achieves state-of-the-art performance among open-source MLLMs, with a 5% performance gain. These results show that training with HAICTrain can significantly improve the caption quality of human actions. Furthermore, using human annotated captions as input can achieve very high accuracy, which means there is still considerable room for improvement. Note that the accuracies of some classes are lower than 25% because we prompted the model not to answer the question if no clues are found in the caption.
Standard Evaluation Setting. In this setting, videos and questions are directly fed into MLLMs for question answering. The results are presented in Table 3. Our post-trained LLaVA-Video-ActionPro-7B outperforms LLaVA-Video by 2.1%, achieving state-of-the-art performance among open-source MLLMs. On the action-related subtasks, our model surpasses the baseline LLaVA-Video-7B model by 1.3%-5.3%. These findings highlight the effectiveness of our HAICTrain dataset for enhancing human action understanding. Additionally, Gemini significantly outperforms GPT-4o, which is the reason why we leverage it to generate the captions of our HAICTrain dataset.
4.3 Effectiveness of Better Captions
In the training set, we combine our HAICTrain with 200K instruction pairs from LLaVA-Video-178K Zhang et al. (2024b) to prevent catastrophic forgotten. To assess the contribution of 126K video-caption pairs in HAICTrain, we replace them with another 126K samples randomly selected from LLaVA-Video-178K and execute post-training on two different MLLMs. The results are presented in Table 4. With the same amount of data, training with HAICTrain consistently yields better improvements across various action question-answering benchmarks, including MVBench, PerceptionTest, ActivityNet-QA and HAICBench. We also try the same videos in HAICTain with their original captions in WebVid and find using these captions can not improve accuracy. All of these demonstrate that HAICTrain is better for the model to understand human actions.
4.4 Qualitative Analysis
Figure 4: A video caption example in HAICBench.
We conduct a qualitative analysis of our post-trained model on HAICBench, as shown in Figure 4. The video features two human boxers fighting on the stage. As outlined in the figure, before post-training, the model can only provide a coarse description of their actions. However, after post-training with HAICTrain, our model delivers detailed descriptions of the fighters, including their hair color, clothing, and positions in the cage. Besides, it presents a clear sequence of actions, describing how the attack and defense progress over time. This also demonstrates the effectiveness of our caption format. By distinguishing the fighters based on detailed attributes, the caption facilitates better individual recognition, enhancing the viewer’s ability to follow the action.
4.5 Effectiveness for Text-video Generation
To evaluate our method’s effectiveness in text-to-video generation, we use the MovieGenBench dataset Polyak et al. (2024), consisting of 1,003 videos. We first take LLaVA-Video-7B and LLaVA-Video-ActionPro-7B to generate captions, which are then fed to HunyuanVideo Kong et al. and Wanx2.1 Aliyun (2023) to generate videos. Then, five human annotators assess the semantic relevance between videos generated by two captions, classifying them as “Good”, “Same”, or “Bad”. The annotators are asked to make decisions upon the original videos in MovieGen. We measure the GSB score, i.e., the percentage of “Good” and “Same” relative to that of “Bad” and “Same”. Results show that LLaVA-Video-ActionPro-7B outperforms LLaVA-Video-7B in generating captions that lead to more semantically accurate videos. On HunyuanVideo, LLaVA-Video-ActionPro-7B achieves a GSB score of 2.15, and on Wanx2.1, 6.81. LLaVA-Video-7B often produced verbose but less accurate captions, failing to capture the original video’s core actions. These findings highlight our model’s superior ability to comprehend the actions in videos and relay captions that enable high-quality, semantically faithful text-to-video generation. More cases are presented in Appendix E.
5 Conclusion
This study addresses the challenge of human action understanding by introducing a novel two-stage data annotation pipeline, combining data accumulation, human-machine annotation, and human verification. This process produces two datasets: HAICTrain, which significantly enhances human action understanding and generation, and HAICBench, a comprehensive benchmark. Both datasets will be made open-source to advance research and applications in this field, promoting broader impact across human behavior understanding and generation.
6 Limitations
Although our work makes notable progress in human action understanding, it still has several limitations. Despite their comprehensive nature, our HAIC datasets may not cover the entire spectrum of human actions, particularly those involving complex interactions or cultural nuances. Furthermore, our work primarily focuses on visual and textual data, lacking integration with audio data, which could provide additional context for understanding. Future work should aim to incorporate audio and further refine the annotation process to address these limitations.
Appendix A Automatic Annotation
Figure 5: Prompt for Gemini-Pro Re-captioning.
In Section 3.2, we utilize Gemini-1.5-Pro to generate captions for HAICTrain in our defined standardized format. The prompt is detailed in Figure 5.
Appendix B Evaluation datasets
In Section 4.1, we use action subsets from MVBench and PerceptionTest as our benchmarks. Here, we explain our subset selection process in detail.
For MVBench, we choose seven categories whose names include "action": Action Antonym, Action Count, Action Sequence, Action Prediction, Action Localization, Fine-grained Action, and Unexpected Action.
For PerceptionTest, we select questions tagged with any of the five labels related to human actions: Action counting, Action recognition, Adversarial action, Distractor action, and Occlusion (Occluded interactions).
Appendix C QA Pair Generation
Figure 6: Prompt for action interaction QA generation.
Figure 7: Prompt for action details QA generation.
Figure 8: Prompt for action sequence QA generation.
Figure 9: Prompt for action count QA generation.
Figure 10: Prompt for human attribute QA generation.
Figure 11: Prompt caption evaluation setting.
As briefly discussed in Section 3.3, we utilized a large language model (LLM) to assist in generating QA pairs for HAICBench. The prompts used are as follows: for action interaction QA, see Figure 6; for action detail QA, see Figure 7; for action sequence QA, see Figure 8; for action count QA, see Figure 9; and for human attribute QA, see Figure 10.
Appendix D Caption Evaluation Setting
In Section 4.1, we outline our caption evaluation setup, as illustrated in Figure 11. Initially, we prompt MLLMs to generate a caption for each video using a specific prompt. We then combine the generated caption with a question to enable Gemini-1.5-Pro to produce answers.
Figure 12: Videos generated by captions from LLaVA-Video and LLaVA-Video-ActionPro of the first sample in MovieGenBench. The main subject in this case is one woman walking along the street. LLaVA-Video-ActionPro provides a more detailed appearance of the woman than LLaVA-Video.
Figure 13: Videos generated by captions from LLaVA-Video and LLaVA-Video-ActionPro of the 17th sample in MovieGenBench. The main subject in this case is one blue animated character. LLaVA-Video incorrectly identifies the main subject.
Figure 14: Videos generated by captions from LLaVA-Video and LLaVA-Video-ActionPro of the 129th sample in MovieGenBench. The main subjects in this case are five cars lined in two rows. LLaVA-Video incorrectly identifies the number of the main subjects.
Appendix E Case Study on Text-to-Video Generation
We used LLaVA-Video and LLaVA-Video-ActionPro to caption the reference videos (generated by MovieGen) in MovieGenBench. Then, we used these captions to generate videos, using the open source HunyuanVideo ***The “infer-steps” parameter is set to 50. and the closed source Wanx2.1 https://tongyi.aliyun.com/wanxiang/videoCreation respectively. In the manual evaluation, we concluded that the captions generated by LLaVA-Video-ActionPro can give a video that is more consistent and reasonable with the original video. Figure 12 is an example of a single human subject (the first sample in the dataset). It can be seen that in the caption, LLaVA-Video lost the description of details, resulting in inconsistent generated content. Figure 13 is an example of a single animated character subject (the 17th sample in the dataset). It can be seen that in the caption, LLaVA-Video incorrectly identified the number of subjects. Figure 14 is an example of multiple subjects (the 129th sample in the dataset). It can be seen that in the caption, LLaVA-Video incorrectly identified the number of subjects and missed the positional relationship. Overall, the introduction of refined and formatted action description data has greatly improved the model’s ability to understand and retell movements. This is of great significance for current mainstream action understanding, anomaly recognition, motion generation, and text-to-video generation.
Appendix F Potential Risks
The source videos of our HAICBench are from YouTube. YouTube inherently contains social biases, and viewing its videos as representative of ’the world’ can perpetuate hegemonic perspectives. The majority of popular YouTubers are men, and the platform’s video practices often reflect gender biases. YouTube also faces issues with hate content, including radical alt-right and ’alt-lite’ material. These issues are exacerbated by the platform’s recommendation algorithm. Even though we downloaded videos independently, filtering them by view count still subjects us to algorithmic influence. The popularity and monetization dynamics on YouTube shape and are shaped by broader cultural trends, affecting the style and content of uploaded videos.